OpenStudy (niahloves):

Help!!! medal and fan Add. 36 + 40 + (–2)

OpenStudy (niahloves):

@texaschic101

OpenStudy (bleuspectre):

36+40=76 76+(-2)=74

OpenStudy (niahloves):

Thanks

OpenStudy (bleuspectre):

No problem

OpenStudy (niahloves):

Subtract. –11 – (–4) @blackfireskull

OpenStudy (jackthegreatest):

wait -11+4 -7

OpenStudy (blackfireskull):

i got -7

OpenStudy (niahloves):

?

OpenStudy (blackfireskull):

-11-(-4)=-7 sooooo

OpenStudy (jackthegreatest):

yes -7 is the answer, i read the problem wrong

OpenStudy (blackfireskull):

@niahloves your answer would be -7